OpenStudy (anonymous):

HELP PLEASE!!! Write the standard form of the equation for the circle that passes through the points (-9, -16), (-9, 32), and (22, 15). Then identify the center and radius. MULTIPLE CHOICE.

OpenStudy (anonymous):

Use this: The equation of a circle : (x-a)2 + (y-b)2 = r2 (a,b) = center r = radius

OpenStudy (anonymous):

a. (x-3)^2 + (y+9)^2 = 25 (-2, 8), 25 b. (x+2)^2 + (y-8)^2 =625 (-2, 8), 25 c. (x+2)^2 + (y-8)^2 = 625 (-3, 9), 5 d. (x-3)^2 + (y+9)^2 = 25 (-3, 9), 5

OpenStudy (anonymous):

@Blank  @navk

OpenStudy (anonymous):

@Tom_Boy_Rebel how do you plug in the equation?

OpenStudy (anonymous):

Plug in the points in the choices to check the answers. The equation which is satisfied by all the three given points is the right one.

OpenStudy (blank ):

Answer is: b. (x+2)^2 + (y-8)^2 =625 (-2, 8), 25

OpenStudy (anonymous):

Ok I get it. Thank you @navk and @Blank 

OpenStudy (anonymous):

Or you can also find the equation from the points which is quite a long method. For that you need to use the general equation x^2 + y^2 + 2gx + 2fy + c = 0

OpenStudy (anonymous):

Thank you :)

OpenStudy (blank ):

ǝɯoɔlǝʍ ǝɹ,noʎ :)

OpenStudy (anonymous):

@Blank  can you help me with one more?

OpenStudy (blank ):

I'll try but I'm really busy.

OpenStudy (anonymous):

Write the standard form of the equation for the circle that passes through the points (30, -2), (-1, -19), and (-18, 12). Then identify the center and radius.

OpenStudy (anonymous):

a. (x+6)^2 + (y-5)^2 = 625 (6, 5), 25 b. (x-6)^2 +(y-5)^2 = 625 (6, 5), 25 c. (x+5)^2 - (y+6)^2 = 650 (3, 5), 15 d. (x+5)^2 + (y+6)^2 = 625 (6, 2), 25

OpenStudy (anonymous):

@Blank 

OpenStudy (blank ):

Just a second...

OpenStudy (blank ):

b. (x-6)^2 +(y-5)^2 = 625 (6, 5), 25

OpenStudy (anonymous):

Thank you so much! @Blank 

OpenStudy (blank ):

ǝɯoɔlǝʍ ǝɹ,noʎ :)